At 3:15, how many degrees there between the two hands of a clock?....
Answer by LinnW(1048)

In relation to 12 , the hour hand is 3.25/12 * 360 degrees = 97.5 degrees
At 3:15 the minute hand is 1/4 of the way around or 1/4 * 360 = 90 degrees
So the hands are (97.5-90) = 7.5 degrees apart.